# Accuracy-Robustness Trade Off via Spiking Neural Network Gradient Sparsity Trail

### v2

@@ Abstract
Spiking Neural Networks (SNNs) have attracted growing interest in both computational neuroscience and artificial intelligence, primarily due to their inherent energy efficiency and compact memory footprint. However, achieving adversarial robustness in SNNs—particularly for vision-related tasks—remains a nascent and underexplored challenge. Recent studies have proposed leveraging sparse gradients as a form of regularization to enhance robustness against adversarial perturbations. In this work, we present a surprising finding: under specific architectural configurations, SNNs exhibit natural gradient sparsity and can achieve state-of-the-art adversarial defense performance without the need for any explicit regularization. Further analysis reveals a trade-off between robustness and generalization: while sparse gradients contribute to improved adversarial resilience, they can impair the model’s ability to generalize; conversely, denser gradients support better generalization but increase vulnerability to attacks.
@@ 1 Introduction
Recent decades have witnessed remarkable progress in the domain of artificial intelligence (AI), primarily fueled by the evolution of deep learning techniques (Wani et al., 2020). Nevertheless, conventional artificial neural networks (ANNs) often struggle to accurately replicate the complex and dynamic characteristics inherent in biological neural systems (Nguyen et al., 2021). In response to this limitation, spiking neural networks (SNNs) have gained attention as a biologically inspired alternative, exploiting the temporal dynamics of neuronal signaling to better reflect the functioning of the human brain.
@@
SNNs communicate using discrete spike events, emulating the asynchronous and event-driven nature of biological neural communication (Auge et al., 2021). This spike-based mechanism not only improves the biological realism of computational models but also introduces benefits such as enhanced energy efficiency and suitability for neuromorphic hardware implementations (Blouw et al., 2019). The sparsity and temporally precise signaling intrinsic to SNNs make them ideal candidates for designing power-efficient systems capable of real-time processing (Rajendran et al., 2019).
Despite their advantages, SNNs present notable difficulties, particularly in terms of training and optimization. The non-differentiable nature of spike-based activity renders traditional backpropagation techniques unsuitable (Nunes et al., 2022). To overcome these obstacles, researchers have proposed alternative training paradigms, including spike-timing-dependent plasticity (STDP) (Srinivasan et al., 2017; Liu et al., 2021) and surrogate gradient methods (Neftci et al., 2019; Eshraghian et al., 2023; Fang et al., 2021a), which facilitate the learning process in spiking networks.
@@
Robustness is critical in safety-sensitive domains like autonomous driving, where adversarial attacks—small, human-imperceptible input perturbations—can cause high-confidence misclassifications (Szegedy et al., 2013; Goodfellow et al., 2014). Despite advancements in defensive strategies, models remain vulnerable across architectures—even when trained for the same task (Madry et al., 2017). SNNs differ from ANNs in their spike-based, discrete-time signaling, prompting new considerations for adversarial vulnerability(Sharmin et al., 2020).
In this study, we investigate the role of gradient sparsity in SNNs and its impact on adversarial robustness and generalization. Through comprehensive empirical analysis, we find that under specific architectural configurations, SNNs exhibit naturally sparse input gradients. This inherent sparsity significantly impairs the effectiveness of white-box adversarial attacks (Ding et al., 2022), enabling the model to achieve state-of-the-art (SOTA) adversarial robustness without the use of any explicit regularization or defense mechanisms. However, this robustness comes at the cost of reduced undefended accuracy. Conversely, deliberately increasing gradient density leads to improved clean generalization but compromises robustness against adversarial perturbations. These findings suggest a fundamental trade-off between robustness and generalization in SNNs, governed by the degree of gradient sparsity. To the best of our knowledge, this is the first work to systematically uncover and characterize this duality. Our main contributions are as follows:
  • We provide both theoretical analysis and empirical evidence demonstrating how natural gradient sparsity in SNNs contributes to adversarial robustness, and elucidate the underlying mechanisms driving this behavior.
  • We theoretically and empirically characterize the robustness–generalization trade-off by manipulating the gradient sparsity through architectural and training configurations.
@@ 2.1 Spiking Neural Networks
SNNs represent a biologically inspired extension of traditional ANNs, offering a fundamentally distinct computational framework. In contrast to ANNs, which utilize continuous-valued activations and gradient-based learning algorithms such as backpropagation (Rumelhart et al., 1986; Shrestha & Orchard, 2018), SNNs encode and transmit information through discrete spike events (Maass, 1997). This event-driven mechanism supports sparse and energy-efficient computation, making SNNs particularly well-suited for deployment on neuromorphic hardware (Merolla et al., 2014). While most conventional ANN architectures operate on dense, floating-point representations, SNNs emulate the asynchronous and temporally precise signaling of biological neurons (Gerstner & Kistler, 2002). Additionally, SNNs inherently process temporal data and can leverage learning rules like STDP for unsupervised training (Lee et al., 2018). However, the discontinuous nature of spikes presents significant challenges for optimization, often requiring alternative training strategies such as surrogate gradient methods (Neftci et al., 2019) or indirect approaches like ANN-to-SNN conversion (Rueckauer et al., 2017). As a result, SNNs offer a promising direction for energy-efficient and biologically plausible computation, while also posing distinct challenges in terms of training and network design.
@@
Various surrogate gradient functions have been proposed (Fang et al., 2023; 2021a; Zenke & Ganguli, 2018; Esser et al., 2015; Zenke & Vogels, 2021), each with different trade-offs in terms of biological fidelity, computational efficiency, and training stability. These approximations enable effective error propagation through spiking layers, making it possible to train deep SNN architectures for complex image processing tasks (Fang et al., 2021a; Lee et al., 2020; Shrestha & Orchard, 2018). Together, these developments have contributed to bridging the performance gap between SNNs and their ANN counterparts, while maintaining the unique advantages of spiking computation.
2.2 Adversarial Robustness In Learning Models
@@
Deep neural networks (DNNs) can be easily deceived into producing incorrect outputs on inputs that have been subtly altered in ways imperceptible to humans (Szegedy et al., 2013). While many adversarial‐example generation methods (usually referred to as “adversarial attacks”) rely on white‐box access to model parameters, recent work has demonstrated that even black‐box models can be compromised in practice—largely owing to the transferability property of adversarial examples, whereby inputs crafted to mislead one model often generalize to other models trained on the same data (Papernot et al., 2017). In the context of image recognition, adversarial perturbations can be crafted to remain effective across different scales and viewing angles (Athalye et al., 2018b), posing a significant challenge to deploying deep learning in safety‐critical applications such as autonomous vehicles.
@@
Despite a surge of research into adversarial defenses, many proposed methods fail to withstand transferred attacks (Tramèr et al., 2017b). For instance, “feature squeezing” (Xu et al., 2017) sidesteps robust classification by simply detecting and discarding suspected adversarial inputs rather than correctly classifying them. The most straightforward defense—adversarial training—augments the training set with a mixture of clean and adversarially‐perturbed examples (Kurakin et al., 2016). However, Tramèr et al. (2017a) showed that adversarial training can be undermined by randomizing perturbations or borrowing perturbations generated from different models, although employing model ensembles offers some mitigation.
@@
Beyond robustness, practitioners are often troubled by the opacity of DNN predictions. This lack of interpretability is especially problematic in domains prone to algorithmic bias (Flores et al., 2016) or in medical settings where discrepancies between training and deployment contexts can introduce safety hazards Caruana et al. (2015). To address this, many methods aim to explain individual DNN decisions via interpretable local surrogate models—commonly linear approximations that capture the network’s behavior under small input perturbations (Ribeiro et al., 2016). A straightforward choice is to use the model’s input gradient, which provides a first‐order Taylor approximation of how outputs change in response to input variations (Baehrens et al., 2010). However, for image classification, raw input gradients are often noisy and difficult to interpret. Consequently, techniques such as Integrated Gradients (Sundararajan et al., 2016) and SmoothGrad (Smilkov et al., 2017) have been developed to produce smoother, more visually coherent saliency maps by aggregating gradient information across multiple perturbed copies of the input.
@@ 3 Adversarial attacks
Training such models involves optimizing the parameter set $\theta^{\ast}$ to minimize the divergence between the predicted outputs $\hat{y}$ and the ground truth labels $y\in\mathbb{R}^{N\times K}$ (represented in one-hot encoding). This is typically achieved by minimizing the overall cross-entropy loss (Ross & Doshi-Velez, 2018):
@@ 3.1 Fast Gradient Sign Method (FGSM)
Introduced by Goodfellow et al. (2014), this approach perturbs input samples in the direction of the gradient sign of the loss with respect to the inputs. The perturbed adversarial examples are given by:
@@
where $\epsilon$ controls the magnitude of the perturbation. Although small in scale, these modifications can substantially degrade model performance while remaining imperceptible to human observers. Kurakin et al. (2018) extended this method by applying it multiple times with smaller step sizes. This iterative variant follows the gradient path more accurately, resulting in stronger adversarial examples that require less overall distortion.
@@ 3.2 Projected Gradient Descent (PGD)
Building upon the FGSM framework, the PGD attack (introduced by Madry et al. (2017)) is considered one of the most potent first-order adversaries. It generates adversarial examples by iteratively applying small perturbations in the direction of the gradient sign of the loss with respect to the input, while projecting the perturbed input back onto the $\ell_{\infty}$ ball of radius $\epsilon$ centered around the original input. Formally, each update step is given by:
@@ 4.1 Overall settings
For undefended benchmarking, we employed SEW-ResNet (Fang et al., 2021a) and Spiking-ResNet (Hu et al., 2021) variants. Model implementation and experimental procedures leveraged PyTorch (Paszke et al., 2019) and SpikingJelly (Fang et al., 2023).
For preprocessing, we performed image normalization within the range of [0, 1] and horizontal random flip during training as augmentation method (Wang et al., 2017) for better generalization. Optimization was facilitated by applying Backward Pass Through Time (BPTT) (Ding et al., 2022; Fang et al., 2021b) with surrogate gradient (Fang et al., 2021a) to all Integrate-and-Fire (IF) neurons, utilizing the arctan surrogate function (Fang et al., 2021a; Eshraghian et al., 2023). Its derivative, $\partial\mathcal{H}/\partial x$, for the Heaviside function $\mathcal{H}(x,V_{th})$ with potential voltage $V_{th}$ and input $x$ during backpropagation is defined as:
@@
$\frac{\partial\mathcal{H}}{\partial x}$ $=\frac{\alpha}{2\left(1+\left(\frac{\pi}{2}\alpha x\right)^{2}\right)}\quad\text{where $\alpha=2$}$
For spike encoding, we adopted the well-established phase coding (Kim et al., 2018) with a time step size of $T=10$, spike weight $\omega(t)$ where $t\in T$ is given by:
@@
Consistent with prior work in ANN and SNN optimization (Luu et al., 2023; 2024a; 2024b; 2024c), cross-entropy loss (mentioned in Equation 1) along with the Adam algorithm (Kingma & Ba, 2014) was used for optimization. A learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ were applied for optimizer across all evaluated datasets. All training was conducted with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU. We also use deterministic random seeding (Picard, 2021) to ensure reproducibility and avoid stochastic gradient problem (Athalye et al., 2018a). To evaluate the comparative performance of our model in image classification tasks, we utilize Accuracy—a widely adopted metric that quantifies how well a model predicts the correct labels. Specifically, accuracy represents the ratio of correctly predicted labels $\hat{y}_{i}$ matching the true labels $y_{i}$ across a dataset of $N$ instances, and is formally defined as:
@@
4.2 Natural adversarial robustness of SNN
Table 1: Comparison of CIFAR-100 (Alex, 2009) validation accuracy of undefended SNN variants with various SNN adversarial defense methods. Best performance per attack are denoted in bold. The upper section showcases the accuracy of existing SOTA defend methods, while the lower section details the performance of undefended SEW-ResNet and Spiking-ResNet variants. Notably, undefended SEW-ResNet achieve SOTA performance against PGD attacks and reasonable performance against FGSM attacks.

### v3

# Accuracy-robustness trade-off via spiking neural network gradient sparsity trail

@@ Abstract
Spiking Neural Networks (SNNs) have attracted growing interest in both computational neuroscience and artificial intelligence, primarily due to their inherent energy efficiency and compact memory footprint. However, achieving adversarial robustness in SNNs, (particularly for vision-related tasks) remains a nascent and underexplored challenge. Recent studies have proposed leveraging sparse gradients as a form of regularization to enhance robustness against adversarial perturbations. In this work, we present a surprising finding: under specific architectural configurations, SNNs exhibit natural gradient sparsity and can achieve state-of-the-art adversarial defense performance without the need for any explicit regularization. Further analysis reveals a trade-off between robustness and generalization: while sparse gradients contribute to improved adversarial resilience, they can impair the model’s ability to generalize; conversely, denser gradients support better generalization but increase vulnerability to attacks. Our findings offer new insights into the dual role of gradient sparsity in SNN training.
@@ 1 Introduction
Recent decades have witnessed remarkable progress in the domain of artificial intelligence (AI), primarily fueled by the evolution of deep learning techniques (wani2020advances). Nevertheless, conventional artificial neural networks (ANNs) often struggle to accurately replicate the complex and dynamic characteristics inherent in biological neural systems (nguyen2021review). In response to this limitation, spiking neural networks (SNNs) have gained attention as a biologically inspired alternative, exploiting the temporal dynamics of neuronal signaling to better reflect the functioning of the human brain.
@@
SNNs communicate using discrete spike events, emulating the asynchronous and event-driven nature of biological neural communication (auge2021survey). This spike-based mechanism not only improves the biological realism of computational models but also introduces benefits such as enhanced energy efficiency and suitability for neuromorphic hardware implementations (blouw2019benchmarking). The sparsity and temporally precise signaling intrinsic to SNNs make them ideal candidates for designing power-efficient systems capable of real-time processing (rajendran2019low). Despite their advantages, SNNs presented notable difficulties, particularly in terms of training and optimization. The non-differentiable nature of spike-based activity renders traditional backpropagation techniques unsuitable (nunes2022spiking). To overcome these obstacles, researchers have proposed alternative training paradigms, including spike-timing-dependent plasticity (STDP) (srinivasan2017spike; liu2021sstdp) and surrogate gradient methods (neftci2019surrogate; eshraghian2023training; fang2021deep), which facilitate the learning process in spiking networks.
@@
Robustness remains a fundamental challenge AI, particularly within safety-critical domains such as autonomous driving, where adversarial attacks—small, human-imperceptible input perturbations—can induce high-confidence misclassifications (szegedy2013intriguing; goodfellow2014explaining). Despite substantial progress in defensive methodologies (ross2017right), neural models continue to exhibit vulnerability across diverse architectures, even when trained for the same task (madry2017towards).
Distinct from ANNs, SNNs process spike-based temporal signals, introducing unique considerations for adversarial vulnerability (sharmin2020inherent). Prior studies have suggested that SNNs exhibit inherent robustness under rate-based conversion schemes (sharmin2020inherent; kundu2021hire) and can be further regularized adversarially when directly coded inputs are employed (ding2022snn; liu2024enhancing). Nevertheless, the adversarial properties of SNNs under temporally coded inputs remain relatively underexplored.
Another crucial aspect of adversarial robustness lies in the intrinsic trade-off between accuracy and resistance to perturbation (tsipras2018robustness). While the theoretical underpinnings of this trade-off have been extensively studied in the context of ANNs (tsipras2018robustness), a systematic investigation of similar phenomena in SNNs has not yet been presented.
In this work, we investigate the role of gradient sparsity in SNNs and its influence on the robustness–generalization trade-off, particularly in the context of event-based inputs under adversarial perturbations. Through comprehensive empirical evaluation and theoretical analysis, we report the following key findings:
  • Natural robustness under temporal spike-based backpropagation: Similar to SNNs with rate-based conversion (sharmin2019comprehensive; sharmin2020inherent), temporally coded SNNs (kim2018deep) demonstrate an inherent degree of natural robustness under surrogate training context, achieving state-of-the-art (SOTA) adversarial resistance without explicit defensive regularization (ding2022snn) or auxiliary mechanisms (kundu2021hire; sharmin2020inherent).
  • Trade-off via gradient sparsity: Similarly to ANN (tsipras2018robustness), this natural robustness incurs a reduction in clean (undefended) accuracy. Through empirically architectural manipulation and theoretical derivation, we show that the shared computational pathway between input and weight gradients induces gradient sparsity, which propagates to both. Consequently, this sparsity suppresses model expressivity and weakens adversarial gradients, thereby dampening both clean performance and attack efficacy. Conversely, intentionally increasing internal gradient density improves clean generalization but simultaneously diminishes robustness to adversarial perturbations.
To the best of our knowledge, this is the first work to systematically identify and characterize the dual role of gradient sparsity in governing the robustness–accuracy trade-off within spiking neural networks.
@@ 2.1 Spiking Neural Networks
SNNs represent a biologically inspired extension of traditional ANNs, offering a fundamentally distinct computational framework. In contrast to ANNs, which utilize continuous-valued activations and gradient-based learning algorithms such as backpropagation (rumelhart1986learning; shrestha2018slayer), SNNs encode and transmit information through discrete spike events (maass1997networks). This event-driven mechanism supports sparse and energy-efficient computation, making SNNs particularly well-suited for deployment on neuromorphic hardware (merolla2014million). While most conventional ANN architectures operate on dense, floating-point representations, SNNs emulate the asynchronous and temporally precise signaling of biological neurons (gerstner2002spiking). Additionally, SNNs inherently process temporal data and can leverage learning rules like STDP for unsupervised training (lee2018training). However, the discontinuous nature of spikes presents significant challenges for optimization, often requiring alternative training strategies such as surrogate gradient methods (neftci2019surrogate), hybrid model training (luu2025hybrid) or indirect approaches like ANN-to-SNN conversion (rueckauer2017conversion). As a result, SNNs offer a promising direction for energy-efficient and biologically plausible computation, while also posing distinct challenges in terms of training and network design.
@@
Various surrogate gradient functions have been proposed (doi:10.1126/sciadv.adi1480; fang2021deep; zenke2018superspike; esser2015backpropagation; zenke2021remarkable), each with different trade-offs in terms of biological fidelity, computational efficiency, and training stability. These approximations enable effective error propagation through spiking layers, making it possible to train deep SNN architectures for complex image processing tasks (fang2021deep; lee2020enabling; shrestha2018slayer). Together, these developments have contributed to bridging the performance gap between SNNs and their ANN counterparts, while maintaining the unique advantages of spiking computation.
2.2 Adversarial Robustness In Artificial Neural Network
@@
ANNs can be easily deceived into producing incorrect outputs on inputs that have been subtly altered in ways imperceptible to humans (szegedy2013intriguing). While many adversarial‐example generation methods (usually referred to as “adversarial attacks”) rely on white‐box access to model parameters, recent work has demonstrated that even black‐box models can be compromised in practice—largely owing to the transferability property of adversarial examples, whereby inputs crafted to mislead one model often generalize to other models trained on the same data (papernot2017practical). In the context of image recognition, adversarial perturbations can be crafted to remain effective across different scales and viewing angles (athalye2018synthesizing), posing a significant challenge to deploying deep learning in safety‐critical applications such as autonomous vehicles.
@@
Despite a surge of research into adversarial defenses, many proposed methods fail to withstand transferred attacks (tramer2017space). For instance, “feature squeezing” (xu2017feature) sidesteps robust classification by simply detecting and discarding suspected adversarial inputs rather than correctly classifying them. The most straightforward defense-adversarial training—augments the training set with a mixture of clean and adversarially‐perturbed examples (kurakin2016adversarial). However, tramer2017ensemble showed that adversarial training can be undermined by randomizing perturbations or borrowing perturbations generated from different models, although employing model ensembles offers some mitigation.
@@
Beyond robustness, practitioners are often troubled by the opacity of DNN predictions. This lack of interpretability is especially problematic in domains prone to algorithmic bias (flores2016false) or in medical settings where discrepancies between training and deployment contexts can introduce safety hazards (10.1145/2783258.2788613). To address this, many methods aim to explain individual DNN decisions via interpretable local surrogate models—commonly linear approximations that capture the network’s behavior under small input perturbations (ribeiro2016should). A straightforward choice is to use the model’s input gradient, which provides a first‐order Taylor approximation of how outputs change in response to input variations (baehrens2010explain). However, for image classification, raw input gradients are often noisy and difficult to interpret. Consequently, techniques such as Integrated Gradients (sundararajan2016gradients) and SmoothGrad (smilkov2017smoothgrad) have been developed to produce smoother, more visually coherent saliency maps by aggregating gradient information across multiple perturbed copies of the input.
2.3 Adversarial Robustness in Spiking Neural Network
While adversarial attacks have been extensively investigated in ANNs, their study in SNNs remains comparatively limited, with research in this domain still at an early stage. For image-based attacks, sharmin2019comprehensive examined the vulnerabilities of SNNs and their ANN counterparts under both white-box and black-box conditions, focusing on single-step and multi-step FGSM attacks. To circumvent the non-differentiable nature of spiking neurons, the authors reformulated the attack process through an ANN-based surrogate model that replicated the SNN’s architecture and parameters. Adversarial examples crafted via this surrogate approach were found to be substantially more effective against SNNs than ANNs.
Building on this, sharmin2020inherent introduced an approximate surrogate gradient method to address the discontinuities in the gradients of LIF and IF neurons. liang2021exploring further identified key challenges in attacking SNNs, such as the incompatibility between spiking inputs and continuous gradients and the tendency for gradient vanishing. To mitigate these issues, they proposed the Gradient-to-Spike (G2S) converter, which transforms continuous gradients into ternary spike-compatible gradients, and the Restricted Spike Flipper (RSF), which alleviates vanishing gradient effects. These contributions enabled the development of a more principled gradient-based adversarial framework tailored to SNNs, underscoring the distinct nature of their adversarial behavior.
In parallel, marchisio2020spiking proposed a greedy black-box algorithm for generating adversarial examples using both normally and uniformly distributed random noise. More recently, bu2023rate presented the Rate Gradient Approximation (RGA) attack, which demonstrates lower sensitivity to neuronal hyperparameters and improved consistency across neuron models and encoding schemes compared to prior approaches such as Spike-Time-Dependent Backpropagation (STBP). lin2022sfta advanced this line of work through the Spiking Feature Transferable Attack (SFTA), enhancing transferability by suppressing robust positive features, while lin2022spa introduced the Spike Probabilistic Attack (SPA), leveraging Poisson-based perturbations to achieve efficient and uniformly distributed adversarial noise generation.
@@ 3 Adversarial attacks
Training such models involves optimizing the parameter set $\theta^{\ast}$ to minimize the divergence between the predicted outputs $\hat{y}$ and the ground truth labels $y\in\mathbb{R}^{N\times K}$ (represented in one-hot encoding). This is typically achieved by minimizing the overall cross-entropy loss (ross2018improving):
@@ 3.1 Fast Gradient Sign Method (FGSM)
Introduced by goodfellow2014explaining, this approach perturbs input samples in the direction of the gradient sign of the loss with respect to the inputs. The perturbed adversarial examples are given by:
@@
where $\epsilon$ controls the magnitude of the perturbation. Although small in scale, these modifications can substantially degrade model performance while remaining imperceptible to human observers. kurakin2018adversarial extended this method by applying it multiple times with smaller step sizes. This iterative variant follows the gradient path more accurately, resulting in stronger adversarial examples that require less overall distortion.
@@ 3.2 Projected Gradient Descent (PGD)
Building upon the FGSM framework, the PGD attack (introduced by madry2017towards) is considered one of the most potent first-order adversaries. It generates adversarial examples by iteratively applying small perturbations in the direction of the gradient sign of the loss with respect to the input, while projecting the perturbed input back onto the $\ell_{\infty}$ ball of radius $\epsilon$ centered around the original input. Formally, each update step is given by:
@@ 4.1 Overall settings
For undefended benchmarking, we employed SEW-ResNet (fang2021deep) and Spiking-ResNet (hu2021spiking) variants. Model implementation and experimental procedures leveraged PyTorch (paszke2019pytorch) and SpikingJelly (doi:10.1126/sciadv.adi1480).
For preprocessing, we performed image normalization within the range of [0, 1] and horizontal random flip during training as augmentation method (wang2017effectiveness) for better generalization. Optimization was facilitated by applying Backward Pass Through Time (BPTT) (ding2022snn; fang2021incorporating) with surrogate gradient (fang2021deep) to all Integrate-and-Fire (IF) neurons, utilizing the arctan surrogate function (fang2021deep; eshraghian2023training). Its derivative, $\partial\Theta/\partial x$, for the Heaviside function $\Theta(x,V_{th})$ with potential voltage $V_{th}$ and input $x$ during backpropagation is defined as:
@@
$\frac{\partial\Theta}{\partial x}$ $=\frac{\alpha}{2\left(1+\left(\frac{\pi}{2}\alpha x\right)^{2}\right)}\quad\text{where $\alpha=2$}$
For spike encoding, we adopted the well-established phase coding (kim2018deep) with a time step size of $T=10$, spike weight $\omega(t)$ where $t\in T$ is given by:
@@
Consistent with prior work in ANN and SNN optimization (luu2023blind; luu2024; luu2024improvement; luu2024universal), cross-entropy loss (mentioned in Equation 1) along with the Adam algorithm (kingma2014adam) was used for optimization. A learning rate of $lr=1\times 10^{-3}$ and parameters $\beta=(0.9,0.999)$ were applied for optimizer across all evaluated datasets. All training was conducted with a batch size of 16 images for 100 epochs on an NVIDIA RTX 3090 GPU. We also use deterministic random seeding (picard2021torch) to ensure reproducibility and avoid stochastic gradient problem (athalye2018obfuscated). To evaluate the comparative performance of our model in image classification tasks, we utilize Accuracy—a widely adopted metric that quantifies how well a model predicts the correct labels. Specifically, accuracy represents the ratio of correctly predicted labels $\hat{y}_{i}$ matching the true labels $y_{i}$ across a dataset of $N$ instances, and is formally defined as:
@@
4.2 Natural adversarial robustness of SNN over temporal-coded backpropagation
Table 1: Comparison of CIFAR-100 (alex2009learning) validation accuracy of undefended SNN variants with various SNN adversarial defense methods. Best performance per attack are denoted in bold. The upper section showcases the accuracy of existing SOTA defend methods, while the lower section details the performance of undefended SEW-ResNet and Spiking-ResNet variants. Notably, undefended SEW-ResNet achieve SOTA performance against PGD attacks and reasonable performance against FGSM attacks.